40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3422     Entry time: Fri Aug 13 18:32:48 2010
Author: steve 
Type: Update 
Category: PEM 
Subject: new legs attached to optical table 

Koji, Rana and Steve

 

The floor preparation continued today. The really messy part of drilling and jack-hammering the concrete were done in controlled manner.

Particle counts peaked at 106K of 0.3 micron and 13 K of 1.0 micron particles/ cfmin at the top of SP table.

Legs were installed with the help of Koji and strongman Rana. Professor Rana was very useful lifting the 85 lbs legs. 

 

Atm1, Drilling holes for ribbed bar reinforcement of concrete

Atm2, Koji and Rana are playing hide and seek while installing legs

Atm3-4, 6 Rigid tripod legs attached by 6x3 of 1/4-20x 7/8 screws. The top plates of the legs were torqued to 108 ft/lbs by Rana

Attachment 1: P1060612.JPG  1.549 MB  | Hide | Hide all
P1060612.JPG
Attachment 2: P1060626.JPG  1.240 MB  | Hide | Hide all
P1060626.JPG
Attachment 3: P1060631.JPG  2.333 MB  | Hide | Hide all
P1060631.JPG
Attachment 4: P1060636.JPG  1.062 MB  | Hide | Hide all
P1060636.JPG
ELOG V3.1.3-